# Distributed MIMO Positioning: Fundamental Limit Analysis and User Tracking Framework Design

### v1

@@ Abstract
This paper presents a comprehensive study on the 3D positioning capabilities of users in distributed multiple-input multiple-output (MIMO) systems. Unlike previous studies that mainly rely on idealized isotropic antenna models, we adopt a polarimetric model that takes advantage of effective aperture distribution functions to characterize realistic antenna patterns, placements, and polarization effects. Based on this model, we analyze the fundamental limits of UE positioning using the Fisher information matrix (FIM) and the position error bound (PEB). The FIM is shown to be expressed as a weighted sum of the information contributions from individual access point (AP)–UE pairs, with each contribution interpreted geometrically across distance, azimuth, and elevation dimensions. The impact of the UE tilt and the spatial distribution of APs on the PEBs is further analyzed. As a further advancement, we propose a complete positioning framework from a UE tracking perspective. By integrating a global probability hypothesis density filter and a PEB-aware AP management strategy, the framework enables accurate tracking while optimizing AP scheduling. Finally, we present a distributed MIMO channel measurement campaign to validate the proposed framework. The results demonstrate a centimeter-level tracking accuracy. In addition, the PEB-aware AP management strategy is shown to maintain robust tracking performance while significantly reducing the number of concurrently active APs, thus lowering the overall system overhead.
@@ II System Model
Consider a multi-user distributed MIMO system in which $K$ spatially distributed APs serve $M$ users. Each AP is equipped with an antenna array consisting of $N$ antenna elements, while each user is equipped with a single antenna. Antenna tilting may occur on the UE side, we denote the antenna tilt angle of the $m$-th UE in the vertical direction with $\beta_{m}$. Without loss of generality, we consider uplink communication. The received signals with bandwidth $B$ are sampled with a frequency spacing of $\Delta B$. We assume that all APs and UEs are synchronized. Given the signal $\mathbf{s}_{k,m}$ transmitted from the $m$-th user, the discrete received signal vector $\mathbf{y}(t_{n_{t}},f_{n_{f}})\in\mathbb{C}^{N\times 1}$ at the $k$-th AP can be expressed as
$\begin{split}\mathbf{y}(t_{n_{t}},f_{n_{f}})=\sum_{l=1}^{L_{k,m}}&\mathbf{C}_{\text{R}}\left(\theta_{k,m}^{(l)},\phi_{k,m}^{(l)}\right)\mathbf{A}_{k,m}^{(l)}\mathbf{r}_{\text{T}}\left[\mathbf{C}_{\text{T}}\right]^{\text{T}}s_{k,m}(f_{n_{f}})\\ &\cdot e^{-j2\pi(f_{n_{f}}\tau_{k,m}^{(l)}+v_{k,m}^{(l)}t_{n_{t}})}+\mathbf{n}_{k,m},\end{split}$ (1)
@@
$\mathbf{Y}_{k,m}=[\mathbf{y}(1,1),...,\mathbf{y}(N_{t},N_{f})]^{\text{T}}=\sum_{l=1}^{L_{k,m}}\mathbf{B}(\mathbf{\Theta}_{l})\cdot\gamma_{l}+\mathbf{n}_{k,m},$ (2)
where $\gamma_{l}=\text{Vec}(\mathbf{A}_{l})=\begin{bmatrix}\alpha_{VV}&\alpha_{VH}&\alpha_{HV}&\alpha_{HH}\end{bmatrix}^{\text{T}}$, and
@@
$\begin{bmatrix}c_{T_{V}}^{\prime}\\ c_{T_{H}}^{\prime}\end{bmatrix}=\begin{bmatrix}c_{T_{V}}\cos\beta_{m}-c_{T_{H}}\sin\beta_{m}\\ c_{T_{V}}\sin\beta_{m}+c_{T_{H}}\cos\beta_{m}\end{bmatrix}$. (4)
The EADF has been widely used to obtain full-directional antenna responses based on measured radiation patterns, which are typically sampled in discrete and limited directions [31]. Specifically, let $\tilde{\mathbf{C}}^{(n)}_{V/H}\in\mathbb{C}^{M_{\theta}\times M_{\phi}}$ denote the measured 3D V/H-polarized radiation pattern of the antenna element $n$, where $M_{\theta}$ and $M_{\phi}$ represent the number of discrete elevation and azimuth angle samples, respectively. Defined as the 2D discrete Fourier transform (2D-DFT) of $\tilde{\mathbf{C}}_{V/H}^{(n)}$, the $(m_{\theta},m_{\phi})$-th element of the EADF $\mathbf{Q}_{V/H}^{(n)}\in\mathbb{C}^{M_{\theta}\times M_{\phi}}$ is computed as $q^{(n)}_{m_{\theta},m_{\phi}}=\bm{\mu}^{\text{T}}_{m_{\theta}}\tilde{\mathbf{C}}^{(n)}_{V/H}\bm{\mu}_{m_{\phi}}$, where $\bm{\mu}_{m_{\theta}}=\exp\left(-j\pi\frac{2(m_{\theta}-1)-M_{\theta}}{M_{\theta}}[0,...,M_{\theta}-1]^{\text{T}}\right),$ and $\bm{\mu}_{m_{\phi}}=\exp\left(-j\pi\frac{2(m_{\phi}-1)-M_{\phi}}{M_{\phi}}[0,...,M_{\phi}-1]^{\text{T}}\right).$ Next, by applying the 2D inverse DFT (IDFT) to $\mathbf{Q}_{V/H}$, the response of the array in an arbitrary direction $(\theta,\phi)$ can be reconstructed as
$\begin{split}\mathbf{c}_{V/H}(\theta,\phi)=&\left[c_{V/H}^{(1)}(\theta,\phi),...,c_{V/H}^{(N)}(\theta,\phi)\right]^{\text{T}}\\ =&\mathbf{Q}_{V/H}\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi})\\ =&\left[\text{Vec}\left(\mathbf{Q}_{V/H}^{(1)}\right)^{\text{T}},...,\text{Vec}\left(\mathbf{Q}_{V/H}^{(N)}\right)^{\text{T}}\right]^{\text{T}}\\ &\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}),\end{split}$ (5)
@@
A key advantage of the above EADF-based antenna model is that it is entirely derived from measured radiation patterns $\tilde{\mathbf{C}}^{(n)}_{V/H}$, thus reflecting real-world characteristics of antenna arrays and inherently capturing realistic effects such as manufacturing imperfections, frame asymmetry, and mutual coupling.
@@ III-A Position Error Bound
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}=\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{&\left(c_{T_{V}}^{\prime}\mathbf{A}^{\theta}_{V}+c_{T_{H}}^{\prime}\mathbf{A}^{\theta}_{H}\right)^{\text{H}}\\ &\cdot\left(c_{T_{V}}^{\prime}\mathbf{A}^{\theta}_{V}+c_{T_{H}}^{\prime}\mathbf{A}^{\theta}_{H}\right)\bigg\},\end{split}$ (12)
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\phi\phi}=\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{&\left(c_{T_{V}}^{\prime}\mathbf{A}^{\phi}_{V}+c_{T_{H}}^{\prime}\mathbf{A}^{\phi}_{H}\right)^{\text{H}}\\ &\cdot\left(c_{T_{V}}^{\prime}\mathbf{A}^{\phi}_{V}+c_{T_{H}}^{\prime}\mathbf{A}^{\phi}_{H}\right)\bigg\},\end{split}$ (13)
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\tau\tau}=\frac{8\pi^{2}E_{s}B_{e}^{2}}{\sigma^{2}}\mathcal{R}\bigg\{&\left(c_{R_{V}}^{\prime}\mathbf{A}^{\tau}_{V}+c_{R_{H}}^{\prime}\mathbf{A}^{\tau}_{H}\right)^{\text{H}}\\ &\cdot\left(c_{R_{V}}^{\prime}\mathbf{A}^{\tau}_{V}+c_{R_{H}}^{\prime}\mathbf{A}^{\tau}_{H}\right)\bigg\},\end{split}$ (14)
where $E_{s}=\sum_{n_{f}=1}^{N_{f}}\|s(f_{n_{f}})\|^{2}$ and $B_{e}=\frac{\sum_{n_{f}=1}^{N_{f}}f_{n_{f}}\|s(f_{n_{f}})\|^{2}}{E_{s}}$ are the transmission signal power and the effective bandwidth [9], respectively. The notations $\mathbf{A}_{V/H}^{\theta}$, $\mathbf{A}_{V/H}^{\phi}$, and $\mathbf{A}_{V/H}^{\tau}$ are expressed as
$\mathbf{A}_{V/H}^{\theta}=\alpha_{VV/HH}\cdot\left(\mathbf{Q}_{R_{V/H}}\cdot\left[(\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta})\otimes\mathbf{w}_{\phi}\right]\right),$ (15)
$\mathbf{A}_{V/H}^{\phi}=\alpha_{VV/HH}\cdot\left(\mathbf{Q}_{R_{V/H}}\cdot\left[\mathbf{w}_{\theta}\otimes(\mathbf{\alpha}_{\phi}\odot\mathbf{w}_{\phi})\right]\right),$ (16)
$\mathbf{A}_{V/H}^{\tau}=\alpha_{VV/HH}\cdot\left(\mathbf{Q}_{R_{V/H}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right).$ (17)
@@
[…]
@@
Assume that the rotation $\mathbf{r}_{k}^{\text{AP}}$ of the AP $k$ causes the x-axis of its LCS to deviate by an angle $\omega_{k}$ counterclockwise with respect to the x-axis of the GCS. The joint FIM of $\mathbf{P}^{\text{UE}}$ is derived as
@@ III-B Impact of Spatial Distribution of APs
In this subsection, the impact of the spatial distribution of APs on the UE position error bounds is analyzed. To simplify the problem, we assume a scenario with azimuth-only signal propagation, where the 2D positions of the APs are given by $\mathbf{p}^{\text{AP}}_{k}=[x^{\text{AP}}_{k},y^{\text{AP}}_{k}]$, and are optimized to minimize the PEB. Consequently, elevation information $\frac{\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}}{c^{2}\tau_{k}^{2}}\mathbf{U}\left(\theta_{k}+\frac{\pi}{2},\phi_{k}\right)$ in (20) is omitted, and (22) is simplified as
@@
where $\tilde{\mathbf{U}}(\phi)\triangleq[\cos\phi\quad\sin\phi]^{\text{T}}\cdot[\cos\phi\quad\sin\phi]$. Clearly, $\mathbf{F}_{\mathbf{p}}$ is affected by both the propagation distance (which is determined by the AP distributions) and the AP rotations. To derive an analytical expression for the optimal AP distribution, we keep the propagation distance and the angle between the UE and the orientation of each AP fixed. This implies that once an AP’s position is determined, its rotation (which is not part of the optimization objective) is also fixed. Under these assumptions, we derive the following theorem.
@@
Given fixed propagation distances and fixed relative angles between the UE and the orientations of the APs, the optimal AP distribution satisfies
@@ III-C Impact of UE Tilt
In this subsection, we evaluate the UE positioning bounds by incorporating the impact of the UE tilt, which is quite common in real scenarios and especially among smartphone users. As defined in Section II, the UE tilt is modeled by a random antenna tilt angle $\beta_{m}$. To derive an analytical PEB expression in terms of $\beta_{m}$, we consider a simplified scenario where the UE is equipped with a single-polarized omnidirectional antenna, i.e., $c_{R_{V}}=1$ while $c_{R_{H}}=0$. Then (4) can be simplified to $\begin{bmatrix}c_{R_{V}}^{\prime}&c_{R_{H}}^{\prime}\end{bmatrix}^{\text{T}}=\begin{bmatrix}\cos\beta_{m}&\sin\beta_{m}\end{bmatrix}^{\text{T}}$. Several distributions have been widely used to model random angles, e.g., normal, Laplacian, and Von Mises distributions. Among them, the Von Mises distribution is adopted here due to its natural suitability for circular random variables [35]. Specifically, we assume that $\beta_{m}$ follows a Von Mises (VM) distribution with
@@
Assume $\beta_{m}$ follows the Von Mises distribution in (28) and the averaged FIM $\bar{\mathbf{F}}_{\mathbf{p}}$ is obtained by replacing $\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}$ in (22) with its expectation $\mathbb{E}\left\{\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}\right\}$. Furthermore, under the assumption that $\left(\mathbf{c}_{V}\right)^{\mathrm{H}}\mathbf{c}_{H}+\left(\mathbf{c}_{H}\right)^{\mathrm{H}}\mathbf{c}_{V}\ll\left(\mathbf{c}_{V}\right)^{\mathrm{H}}\mathbf{c}_{V}+\left(\mathbf{c}_{H}\right)^{\mathrm{H}}\mathbf{c}_{H}$, where $\mathbf{c}_{V/H}$ is defined in (5), $\mathbb{E}\left\{\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}\right\}$ has the following analytical expression:
$\begin{split}\mathbb{E}\left\{\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}\right\}=\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{&\frac{1+\rho}{2}(\mathbf{A}_{V}^{z})^{\mathrm{H}}\mathbf{A}_{V}^{z^{\prime}}\\ &+\frac{1-\rho}{2}(\mathbf{A}_{H}^{z})^{\mathrm{H}}\mathbf{A}_{H}^{z^{\prime}}\bigg\},\end{split}$ (29)
where $\rho=\frac{I_{2}(\kappa_{m})}{I_{0}(\kappa_{m})}\cos(2\mu_{m})$. The terms $\mathbf{A}_{V/H}^{\theta}$, $\mathbf{A}_{V/H}^{\phi}$, and $\mathbf{A}_{V/H}^{\tau}$ have been defined in (15)–(17), while for the other terms,
$\mathbf{A}_{\mathrm{p}}^{\mathcal{R}\{\alpha_{\mathrm{pp}}\}}=\alpha_{\mathrm{pp}}\left(\mathbf{Q}_{R_{\mathrm{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right),$ (30)
$\mathbf{A}_{\mathrm{p}}^{\mathcal{I}\{\alpha_{\mathrm{pp}}\}}=j\alpha_{\mathrm{pp}}\left(\mathbf{Q}_{R_{\mathrm{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right),$ (31)

### v2

@@ Abstract
This paper presents a comprehensive study on the 3D positioning capabilities in distributed multiple-input multiple-output (MIMO) systems. Unlike previous studies that mainly rely on idealized isotropic antenna models, we adopt a polarimetric model that takes advantage of effective aperture distribution functions to characterize realistic antenna patterns, placements, and polarization effects. Based on this model, we analyze the fundamental limits of UE positioning using the Fisher information matrix (FIM) and the position error bound (PEB). The FIM is shown to be expressed as a weighted sum of the information contributions from individual access point (AP)-UE pairs, with each contribution interpreted geometrically across distance, azimuth, and elevation dimensions. The impact of the UE tilt and the spatial distribution of APs on the PEBs is further analyzed. As a further advancement, we propose a complete positioning framework from a UE tracking perspective. By integrating a global probability hypothesis density filter and a PEB-aware AP management strategy, the framework enables accurate tracking while optimizing AP scheduling. Finally, we present a distributed MIMO channel measurement campaign to validate the proposed framework. The results demonstrate a centimeter-level tracking accuracy. In addition, the PEB-aware AP management strategy is shown to maintain robust tracking performance while significantly reducing the number of concurrently active APs, thus lowering the overall system overhead.
@@ II System Model
Consider a multi-user distributed MIMO system in which $K$ spatially distributed APs serve $M$ users. Each AP is equipped with an antenna array consisting of $N$ antenna elements, while each user is equipped with a single antenna. Antenna tilting may occur on the UE side, we denote the antenna tilt angle of the $m$-th UE in the vertical direction with $\beta_{m}$. Without loss of generality, we consider uplink communication. The received signals with bandwidth $B$ are sampled with a frequency spacing of $\Delta B$. We assume that all APs and UEs are synchronized. Given the signal $\mathbf{s}_{m}$ transmitted from the $m$-th user, the discrete received signal vector $\mathbf{y}(t_{n_{t}},f_{n_{f}})\in\mathbb{C}^{N\times 1}$ at the $k$-th AP can be expressed as
$\begin{split}\mathbf{y}(t_{n_{t}},f_{n_{f}})=\sum_{l=1}^{L_{k,m}}&\mathbf{C}_{\text{R}}\left(\theta_{k,m}^{(l)},\phi_{k,m}^{(l)}\right)\mathbf{A}_{k,m}^{(l)}\mathbf{r}_{\text{T}}\left[\mathbf{C}_{\text{T},m}\right]^{\text{T}}\mathbf{s}_{m}(f_{n_{f}})\\ &\cdot e^{-j2\pi(f_{n_{f}}\tau_{k,m}^{(l)}+v_{k,m}^{(l)}t_{n_{t}})}+\mathbf{n}_{k,m},\end{split}$ (1)
@@
$\mathbf{Y}_{k,m}=[\mathbf{y}(t_{1},f_{1}),...,\mathbf{y}(t_{N_{t}},f_{N_{f}})]^{\text{T}}=\sum_{l=1}^{L_{k,m}}\mathbf{B}(\mathbf{\Theta}_{l})\cdot\gamma_{l}+\mathbf{n}_{k,m},$ (2)
where $\gamma_{l}=\text{Vec}(\mathbf{A}_{l})=\begin{bmatrix}\alpha_{\rm{VV}}&\alpha_{\rm{VH}}&\alpha_{\rm{HV}}&\alpha_{\rm{HH}}\end{bmatrix}^{\text{T}}$, and
@@
$\begin{bmatrix}c_{\rm{T_{V}}}^{\prime}\\ c_{\rm{T_{H}}}^{\prime}\end{bmatrix}=\begin{bmatrix}c_{\rm{T_{V}}}\cos\beta_{m}-c_{\rm{T_{H}}}\sin\beta_{m}\\ c_{\rm{T_{V}}}\sin\beta_{m}+c_{\rm{T_{H}}}\cos\beta_{m}\end{bmatrix}$. (4)
The EADF has been widely used to obtain full-directional antenna responses based on measured radiation patterns, which are typically sampled in discrete and limited directions [31]. Specifically, let $\tilde{\mathbf{C}}^{(n)}_{\rm{V/H}}\in\mathbb{C}^{M_{\theta}\times M_{\phi}}$ denote the measured 3D V/H-polarized radiation pattern of the antenna element $n$, where $M_{\theta}$ and $M_{\phi}$ represent the numbers of discrete elevation and azimuth angle samples, respectively. The rows and columns of $\tilde{\mathbf{C}}^{(n)}_{\rm{V/H}}$ correspond to $0,\frac{2\pi}{M_{\theta}},...,\frac{2\pi(M_{\theta}-1)}{M_{\theta}}]^{\text{T}}$ and $0,\frac{2\pi}{M_{\phi}},...,\frac{2\pi(M_{\phi}-1)}{M_{\phi}}]^{\text{T}}$, respectively. Defined as the 2D discrete Fourier transform (2D-DFT) of $\tilde{\mathbf{C}}_{\rm{V/H}}^{(n)}$, the $(m_{\theta},m_{\phi})$-th element of the EADF $\mathbf{Q}_{V/H}^{(n)}\in\mathbb{C}^{M_{\theta}\times M_{\phi}}$ is computed as $q^{(n)}_{m_{\theta},m_{\phi}}=\bm{\mu}^{\text{T}}_{m_{\theta}}\tilde{\mathbf{C}}^{(n)}_{\rm{V/H}}\bm{\mu}_{m_{\phi}}$, where $\bm{\mu}_{m_{\theta}}=\exp\left(-j\pi\frac{2(m_{\theta}-1)-M_{\theta}}{M_{\theta}}[0,...,M_{\theta}-1]^{\text{T}}\right),$ and $\bm{\mu}_{m_{\phi}}=\exp\left(-j\pi\frac{2(m_{\phi}-1)-M_{\phi}}{M_{\phi}}[0,...,M_{\phi}-1]^{\text{T}}\right).$ Next, by applying the 2D inverse DFT (IDFT) to $\mathbf{Q}_{V/H}$, the response of the array in an arbitrary direction $(\theta,\phi)$ can be reconstructed as
$\begin{split}\mathbf{c}_{\rm{V/H}}(\theta,\phi)=&\left[c_{\rm{V/H}}^{(1)}(\theta,\phi),...,c_{\rm{V/H}}^{(N)}(\theta,\phi)\right]^{\text{T}}\\ =&\mathbf{Q}_{\rm{V/H}}\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi})\\ =&\left[\text{vec}\left(\mathbf{Q}_{\rm{V/H}}^{(1)}\right)^{\text{T}},...,\text{vec}\left(\mathbf{Q}_{\rm{V/H}}^{(N)}\right)^{\text{T}}\right]^{\text{T}}\\ &\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}),\end{split}$ (5)
@@
A key advantage of the above EADF-based antenna model is that it is entirely derived from measured radiation patterns $\tilde{\mathbf{C}}^{(n)}_{\rm{V/H}}$, thus reflecting real-world characteristics of antenna arrays and inherently capturing realistic effects such as manufacturing imperfections, frame asymmetry, and mutual coupling.
@@ III-A Position Error Bound
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}=\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{&\left(c_{\rm{T_{V}}}^{\prime}\mathbf{A}^{\theta}_{\rm{V}}+c_{\rm{T_{H}}}^{\prime}\mathbf{A}^{\theta}_{\rm{H}}\right)^{\text{H}}\\ &\cdot\left(c_{\rm{T_{V}}}^{\prime}\mathbf{A}^{\theta}_{\rm{V}}+c_{\rm{T_{H}}}^{\prime}\mathbf{A}^{\theta}_{\rm{H}}\right)\bigg\},\end{split}$ (12)
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\phi\phi}=\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{&\left(c_{\rm{T_{V}}}^{\prime}\mathbf{A}^{\phi}_{\rm{V}}+c_{\rm{T_{H}}}^{\prime}\mathbf{A}^{\phi}_{\rm{H}}\right)^{\text{H}}\\ &\cdot\left(c_{\rm{T_{V}}}^{\prime}\mathbf{A}^{\phi}_{\rm{V}}+c_{\rm{T_{H}}}^{\prime}\mathbf{A}^{\phi}_{\rm{H}}\right)\bigg\},\end{split}$ (13)
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\tau\tau}=\frac{8\pi^{2}E_{s}B_{e}^{2}}{\sigma^{2}}\mathcal{R}\bigg\{&\left(c_{\rm{T_{V}}}^{\prime}\mathbf{A}^{\tau}_{\rm{V}}+c_{\rm{T_{H}}}^{\prime}\mathbf{A}^{\tau}_{\rm{H}}\right)^{\text{H}}\\ &\cdot\left(c_{\rm{T_{V}}}^{\prime}\mathbf{A}^{\tau}_{\rm{V}}+c_{\rm{T_{H}}}^{\prime}\mathbf{A}^{\tau}_{\rm{H}}\right)\bigg\},\end{split}$ (14)
where $E_{s}=\sum_{n_{f}=1}^{N_{f}}\|s(f_{n_{f}})\|^{2}$ and $B_{e}=\frac{\sum_{n_{f}=1}^{N_{f}}f_{n_{f}}\|s(f_{n_{f}})\|^{2}}{E_{s}}$ are the transmission signal power and the effective bandwidth [9], respectively. The notations $\mathbf{A}_{\rm{V/H}}^{\theta}$, $\mathbf{A}_{\rm{V/H}}^{\phi}$, and $\mathbf{A}_{\rm{V/H}}^{\tau}$ are expressed as
$\mathbf{A}_{\rm{V/H}}^{\theta}=\alpha_{\rm{VV/HH}}\cdot\left(\mathbf{Q}_{\rm{R_{V/H}}}\cdot\left[(\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta})\otimes\mathbf{w}_{\phi}\right]\right),$ (15)
$\mathbf{A}_{\rm{V/H}}^{\phi}=\alpha_{\rm{VV/HH}}\cdot\left(\mathbf{Q}_{\rm{R_{V/H}}}\cdot\left[\mathbf{w}_{\theta}\otimes(\mathbf{\alpha}_{\phi}\odot\mathbf{w}_{\phi})\right]\right),$ (16)
$\mathbf{A}_{\rm{V/H}}^{\tau}=\alpha_{\rm{VV/HH}}\cdot\left(\mathbf{Q}_{\rm{R_{V/H}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right).$ (17)
@@
As only the parameters $\theta_{k}$, $\phi_{k}$, and $\tau_{k}$ are of interest to determine $\mathbf{p}^{\text{UE}}$, we adopt the notion of equivalent FIM (EFIM) of $\mathbf{F}_{\xi_{k}}$ [10], defined as $\mathbf{F}_{\bm{\xi}_{k}}=\mathbf{F}_{\xi\xi}-\mathbf{F}_{\xi\alpha}\mathbf{F}_{\alpha\alpha}\mathbf{F}_{\xi\alpha}^{\text{T}}$. The FIM of $\mathbf{p}^{\text{UE}}$ from the $k$-th AP, under its LCS, can then be obtained using the chain rule [33] as
@@
Assume that the orientation $\mathbf{r}_{k}^{\text{AP}}$ of the AP $k$ causes the x-axis of its LCS to deviate by an angle $\omega_{k}$ counterclockwise with respect to the x-axis of the GCS. The joint FIM of $\mathbf{P}^{\text{UE}}$ is derived as
@@ III-B Impact of Spatial Distribution of APs
In this subsection, the impact of the spatial distribution of APs on the UE position error bounds is analyzed. To simplify the problem, we assume a scenario with azimuth-only signal propagation, where the 2D positions of the APs are given by $\mathbf{p}^{\text{AP}}_{k}=[x^{\text{AP}}_{k},y^{\text{AP}}_{k}]$. Consequently, elevation information $\frac{\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}}{c^{2}\tau_{k}^{2}}\mathbf{U}\left(\theta_{k}+\frac{\pi}{2},\phi_{k}\right)$ in (20) is omitted, and (22) is simplified as
@@
where $\tilde{\mathbf{U}}(\phi)\triangleq[\cos\phi\quad\sin\phi]^{\text{T}}\cdot[\cos\phi\quad\sin\phi]$.
Our target is to minimize the PEB through optimizing AP distributions $\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{K}]$. Clearly, $\mathbf{F}_{\mathbf{p}}$ is affected by both the propagation distance (which is determined by the AP distributions) and the AP orientations. To derive an analytical expression for the optimal AP distribution, we keep the propagation distance and the angle between the UE and the orientation of each AP fixed. This implies that once an AP’s position is determined, $w_{k}$ (which is not part of the optimization objective) is also fixed. Under these assumptions, we derive the following theorem.
@@
Given fixed propagation distances and fixed orientations of the APs, the optimal AP distribution satisfies
@@ III-C Impact of UE Tilt
In this subsection, we evaluate the UE positioning bounds by incorporating the impact of the UE tilt, which is quite common in real scenarios and especially among smartphone users. As defined in Section II, the UE tilt is modeled by a random antenna tilt angle $\beta_{m}$. To derive an analytical PEB expression in terms of $\beta_{m}$, we consider a simplified scenario where the UE is equipped with a single-polarized omnidirectional antenna, i.e., $c_{\rm{R_{V}}}=1$ while $c_{\rm{R_{H}}}=0$. Then (4) can be simplified to $\begin{bmatrix}c_{\rm{R_{V}}}^{\prime}&c_{\rm{R_{H}}}^{\prime}\end{bmatrix}^{\text{T}}=\begin{bmatrix}\cos\beta_{m}&\sin\beta_{m}\end{bmatrix}^{\text{T}}$. Several distributions have been widely used to model random angles, e.g., normal, Laplacian, and Von Mises distributions. Among them, the Von Mises distribution is adopted here due to its natural suitability for circular random variables [35]. Specifically, we assume that $\beta_{m}$ follows a Von Mises (VM) distribution with
@@
Assume $\beta_{m}$ follows the Von Mises distribution in (28) and the averaged FIM $\bar{\mathbf{F}}_{\mathbf{p}}$ is obtained by replacing $\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}$ in (22) with its expectation $\mathbb{E}\left\{\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}\right\}$. Furthermore, under the assumption that $\left(\mathbf{c}_{\rm{V}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{H}}+\left(\mathbf{c}_{\rm{H}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{V}}\ll\left(\mathbf{c}_{\rm{V}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{V}}+\left(\mathbf{c}_{\rm{H}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{H}}$, where $\mathbf{c}_{\rm{V/H}}$ is defined in (5), $\mathbb{E}\left\{\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}\right\}$ has the following analytical expression:
$\begin{split}\mathbb{E}\left\{\left[\mathbf{F}_{\Theta_{k}}\right]_{zz^{\prime}}\right\}=\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{&\frac{1+\rho}{2}(\mathbf{A}_{\rm{V}}^{z})^{\mathrm{H}}\mathbf{A}_{\rm{V}}^{z^{\prime}}\\ &+\frac{1-\rho}{2}(\mathbf{A}_{\rm{H}}^{z})^{\mathrm{H}}\mathbf{A}_{\rm{H}}^{z^{\prime}}\bigg\},\end{split}$ (29)
where $\rho=\frac{I_{2}(\kappa_{m})}{I_{0}(\kappa_{m})}\cos(2\mu_{m})$. The terms $\mathbf{A}_{\rm{V/H}}^{\theta}$, $\mathbf{A}_{\rm{V/H}}^{\phi}$, and $\mathbf{A}_{\rm{V/H}}^{\tau}$ have been defined in (15)–(17), while for the other terms,
$\mathbf{A}_{\mathrm{p}}^{\mathcal{R}\{\alpha_{\mathrm{pp}}\}}=\alpha_{\mathrm{pp}}\left(\mathbf{Q}_{\rm{R}_{\mathrm{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right),$ (30)
$\mathbf{A}_{\mathrm{p}}^{\mathcal{I}\{\alpha_{\mathrm{pp}}\}}=j\alpha_{\mathrm{pp}}\left(\mathbf{Q}_{\rm{R}_{\mathrm{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right),$ (31)